OpenStudy (slickster256):

Which of the following is the solution to the equation 4 over 5n = 20? n = 4 n = 12 n = 16 n = 25

OpenStudy (igreen):

Divide 5 to both sides, what do you get?

OpenStudy (slickster256):

4

OpenStudy (igreen):

\(5n = 20\) \(\dfrac{5n}{5} = \dfrac{20}{5}\) \(\dfrac{\cancel 5n}{\cancel 5} = \dfrac{20}{5}\) \(n = \dfrac{20}{5}\) \(n = 4\) \(\Huge\color{lime} \checkmark\)

OpenStudy (welshfella):